(3x/5)+(1/7)=(2x/7)-3
1
Expert's answer
2012-06-08T07:51:14-0400
(3x*7+1*5)/(7*5)=(2x-3*7)/7
21x+5=5*(2x-21)
21x+5=10x-105
11x=-110
x=-10